Consider a cash flow and interest profile as shown:

Year 0 Year 1 Year 2 Year 3
Cash Flow at the end of the year -1000 3700 2800 1200
Interest Rate that year N.A. 4% 7%

11.5%

The worth at the end of Year 3 of these cash flows is:
Answer #1

F = P *(1+I)^t

Future worth = -1000 * (1+0.04) *(1+0.07) *(1+0.115) + 3700 *(1+0.07) *(1+0.115) + 2800 *(1+0.115) + 1200

= -1000 * (1.04) *(1.07) *(1.115) + 3700 *(1.07) *(1.115) + 2800 *(1.115) + 1200

= 7495.51